Curriculum Vitae

Shardha Jogee

Address Contact Dr. Shardha Jogee Work: (626) 395 – 2121 Division of Physics, Mathematics, and Astronomy Home: (626) 440 –9453 California Institute of Technology Fax: (626) 568 –9352 1200 E. California Blvd., MS 105-24 E-mail: sj@astro.caltech.edu Pasadena , CA 91125 U.S.A. http://www.astro.caltech.edu/∼sj/

Date of Birth: April 7, 1970 Nationality: Mauritian

Education • Ph.D., Astronomy, Yale University, 1999

• M.A., Physics, Cambridge University, England, 1995

• M.Phil., Astronomy, Yale University, 1994

• M.S., Astronomy, Yale University, 1994

• B.A. Honors in Physics, Cambridge University, England, 1992

Research Interests • The Molecular ISM in Starbursts, AGN, and Ultra Luminous IR galaxies

• Bar and Gas Dynamics: Implications for Galaxy Evolution

• Fueling and Evolution of Nuclear Starbursts and AGN

• Star Formation and Galaxy Interactions

Positions/Experience • 1999 – Present: Postdoctoral Scholar, California Institute of Technology (Caltech)

• 2000: Participant in VLA Summer School on Radio Interferometry, New Mexico

• 1997: Summer Research Assistant, Yale University:

‘ The Molecular Gas Dynamics and Starburst-Driven Outflow in NGC 2782’

• 1996: Participant in the Summer School at Les Houches Ecole de Physique, France • 1993: Summer Research Assistant, Yale University.

‘Implementing a new formulation of convection theory for stellar evolution’

• 1994 – 99: Candidate in the Ph.D. program in Astronomy, Yale University

• 1992 – 94: Candidate in the M.S. and M.Phil program, Yale University

• 1989 – 1992: Undegraduate student, Cambridge University, England Research project: ‘Simulations of a Percolation Model for Star Formation in Spirals’

Responsibilities/Experience • 2000 – Present: Referee for articles in ApJ, ApJL, A&A, and conference proceedings.

• 1999 – Present: Member of the Time Allocation Committee for the Owens Valley Radio Observatory (OVRO)

• 1999 – Present: Responsibilities in the OVRO group include: (1) Processing proposals submitted to the OVRO mm array and coordinating logistics of the Time Allocation Committee, (2) Editor of the first observing manual for the OVRO mm array, (3) Flux calibration for the OVRO database, (4) Technical support for external observers,

• 1999 – 2000: Organizer of the Caltech Astronomy Department Weekly Science Talks • 1996 – 97: Organizer of a seminar series on ‘Extragalactic Astronomy’, Yale University • 1995 – 98: Coordinator for the Yale University Graduate Housing

Observing/Data Analysis Experience • Owens Valley Radio Observatory (mm interferometry)

• Palomar 5-m and 1.5-m (imaging and spectroscopy)

• Keck II 10-m (echelle spectroscopy)

• Kitt Peak National Observatory 0.9-m, 4.0-m (imaging)

• Wisconsin Indiana Yale NOAO 3.5-m (imaging)

• Isaac Newton 2-5-m, William Herschel 4.2-m (imaging)

• Canada-France-Hawaii 3.6-m Telecope (imaging)

• Hubble Space Telescope (archival images)

• VLA (archival 21 cm HI line and 20 cm radio continuum data)

• Routinely use data reduction software such as AIPS, IDL, IRAF, ZODIAC, KARMA • Programming experience in Fortran, IDL, Meta Language (ML), and Pascal

• Experience in multi-wavelength (mm, near-infrared, optical and X-ray) data analysis and close collaborations with theorists (e.g., Profs. R. Larson and I. Shlosman) for modeling of data and constraining theory

Teaching Experience

• 2001: Guest lecturer on ‘Gas Dynamics and ISM’ for graduate/undegraduate course, Caltech

• 1999 – Present: Supervision of two graduate students for research projects, Caltech • 1999 – Present: Training of new OVRO observers in array operations and data reduction • 1994 – 96: Lead teaching assistant, guest lecturer, and discussion leader for undergrad-

uate astronomy courses, Yale University

• 1992 – 93: Junior teaching assistant for undergraduate astronomy courses, Yale University

Oral Presentations

• July 2001: Center for Star Formation Studies Workshop, Santa Cruz, ed. D. Hollenbach: ‘Starburst: Triggers and Evolution’ (invited contribution)

• May 2001: La Palma Conference on The Central kpc of Starbursts and AGN: ‘Molecular Gas in The Central Kpc of Starbursts and AGN’

• Sep. 2000: Rinberg Workshop, Starbursts Near and Far, eds L. Tacconi & D. Lutz (Springer): ‘Circumnuclear Gas Transport and Star Formation Triggers’

• Aug. 2000: IAU Symposium 205, Galaxies and their Constituents at the Highest Angular Resolution: ‘Probing Circumnuclear Star Formation Efficiency at High Resolution’

• 1999 – Present: Multiple seminars at Caltech

• Language skills: Fluency in French and English. Knowledge of Spanish Awards:

• American Association of University Women Educational Fellowship (AAUWEF), 1996 • Amelia Earheart Fellowship, Zonta International, 1996

• Yale University J. F. Enders Fellowship, Yale University, 1995

• Sigma Xi Grants-in-Aid of Research, Sigma Xi Society, 1995

• Yale University J. F. Enders Research Grant, Yale University, 1995

• Garfinkel Prize, Yale University, 1992

• Elected to status of Fellow, Cambridge University, England, 1990-1992

• Full undergraduate scholarship in Physics, Cambridge University, England, 1989-1992